You are required to construct 90% confidence interval for mean life time of tyres from a sample of 250 tyres whose mean is found to be 35,290kms. The population standard deviation is 1024 kms.


1
Expert's answer
2021-02-10T02:37:00-0500

"n = 250"

"\bar x = 35290"

"\sigma =1024"

The 90% confidence interval for mean is given by:

"\bar x\pm Z_{\frac{\alpha}{2}}\frac{\sigma}{\sqrt n}=\bar x\pm Z_{0.05}\frac{\sigma}{\sqrt n}=35290\pm1.64\cdot\frac{1024}{\sqrt{250}}=35290\pm106"

So the 90% confidence interval estimate for the mean life time of tyres is (35184, 35396)